# SOMA: a novel sampler for exchangeable variables

### v1

@@ 5 Numerical experiments
In this section, we empirically compare SOMA with Sys-IMwG and Ran-IMwG samplers, in terms of convergence rates and acceptance probabilities, which have been studied in Section 4. We study a two-component synthetic example in Section 5.1, and data imputation based on perturbed histograms in Section 5.2. Moreover, in Section 5.3, we embed SOMA into a data augmentation MCMC framework for private Bayesian linear regression, highlighting SOMA’s versatility.
5.1 A two-component synthetic example
We construct a $n=2$ permutation invariant distribution under Examples 2, 4 and 5. Let $X_{1},X_{2}\sim\mathrm{Beta}(a_{0},b_{0})$ independently a priori and let $Y\sim\mathrm{Laplace}\left((X_{1}+X_{2})/2,1/\epsilon\right)$. Releasing $Y=y$ instead of the confidential data $X$ satisfied $\epsilon-$DP, and the joint distribution with density
$\pi(x;y,\epsilon)\propto\exp\left(-\frac{|y-x_{1}/2-x_{2}/2|}{2\epsilon}\right)(x_{1}x_{2})^{a_{0}-{}1}(1-x_{1})^{b_{0}-1}(1-x_{2})^{b_{0}-1}$ (25)
is a permutation invariant target density. In this case, the bounded ratio of weights assumption holds for $M=\exp(\epsilon)$.
@@ 5.1 A two-component synthetic example
Figure 2 presents the 2-dimensional trace plots of the three methods over 500 iterations, starting from the same initial state $(0.3,0.3)$ with $a_{0}=b_{0}=10$ and $\epsilon=20$. SOMA traverses the higher density regions of $\pi$ more rapidly and can explore the state space more efficiently. In contrast, Ran-IMwG gets stuck in some local regions, resulting in under exploration of the entire state space.
@@
Figure 2: Trace plots of different methods. The starting point is marked in deep purple, with subsequent samples in lighter colors. The orange density curve represents the target marginal posterior distribution. SOMA achieves better exploration compared to other methods.
Next we quantitatively compare the methods using their empirical acceptance probabilities. In addition to chains targeting the density (25) in Figure 3, we add another target distribution similar to (25), but with $X_{1},X_{2}\sim\mathrm{Exp}(1)$ a priori instead. The target distributions considered in Figure 3-B no longer satisfy Assumption 2. However, SOMA still has a higher acceptance rate than the other two samplers, agreeing with the results proved in Section 4. Also, the empirical acceptance rate of SOMA is better than our theoretical lower bound in (9). This suggests the possibility of improving the convergence analysis of SOMA for specific target and proposal distributions.
@@
Figure 3: Average acceptance rate comparison for varying values of $\epsilon$ in (25). A. Target distribution has $X_{1},X_{2}\sim\mathrm{Beta}(10,10)$ a priori. B. Target distribution has $X_{1},X_{2}\sim\mathrm{Exp}(1)$ a priori. SOMA has higher acceptance rates across all experiment settings.
Using coupled chains, we can estimate the convergence rates using the empirical distribution of the coupling time. This enables us to investigate the impact of posterior correlation $\rho_{\pi}(X_{1},X_{2})$ on convergence speed, illustrated in Figure 4. Higher posterior correlation typically indicates a more challenging sampling scenario, and indeed all three samplers mix more slowly under stronger posterior correlation. In both simulation setups (with or without Assumption 2), SOMA is the most efficient in the high correlation scenario ($\rho<-0.4$) and Sys-IMwG is the most efficient when $\rho$ is small.
A
B
Figure 4: Convergence rates with different posterior correlations. A. Using $\mathrm{Beta}(10,10)$ priors on $X_{i}.$ This setting satisfies $\epsilon$-DP and Assumption 2. B. Using $\mathrm{Exp}(1)$ priors. The bounded ratio assumption is violated.
5.2 Data imputation from perturbed histograms
We apply SOMA to data imputation from perturbed histograms (Wasserman and Zhou,, 2010). This is a specific example of Example 2: data augmentation after differential privacy. Suppose $X_{i}$ are iid from some distribution $q$ with support $\mathbb{X}$. The perturbed histogram mechanism can release the empirical distribution of the observations while satisfying $\epsilon$-DP. After counting the number of observations in each bin (denoted $C_{j}$ for $1\leq j\leq m$), this mechanism adds independent Laplace noise to the raw counts, resulting in $D_{j}\mid C_{j}\sim\mathrm{Laplace}(C_{j},2/\epsilon)$. We aim to sample from the posterior distribution $X\mid D$.
Our experiments use $q(x)=1$ for $x\in\mathbb{X}=(0,1)$ and privacy level $\epsilon=5$, and $m=10$ equidistant bins. We vary the number of observations from $n=2$ to $n=60$.
A
B
To compare the three samplers, we first compare the traceplot of the quartiles of $X$ given $D$ for $n=20$ and $n=50$. These traceplots suggest that SOMA converges more rapidly.
A
B
A
B
5.3 Bayesian linear regression given privatized data
In this section, we apply SOMA as the data imputation step within a data augmentation MCMC (DAMCMC) algorithm (Van Dyk and Meng,, 2001; Hobert,, 2011; Roy et al.,, 2024). We specifically consider the task of differentially private Bayesian linear regression and use the algorithm in Ju et al., (2022) for baseline comparisons.
In traditional linear regression, there are $n$ observations with $p$ predictors/covariates and one response. With $X\in\mathbb{X}^{n\times p}$ and coefficients $\beta_{0}\in\mathbb{R},\beta\in\mathbb{R}^{p}$, the linear model $Y\sim\mathcal{N}(X\beta+\beta_{0},\sigma^{2})$ assumes homogeneous Gaussian noise. In differentially private linear regression, we observe ‘confidential data’ $x$ and $y$ indirectly through $s_{\mathrm{dp}}$ after adding privacy noise to summary statistics reflecting $x^{\top}y,x^{\top}x,y^{\top}y.$ Privatization also includes a clamping step to that truncates confidential $x,y$ values to a bounded interval. One can refer to Bernstein and Sheldon, (2019); Ju et al., (2022) or Section C.2 for more details on this privacy mechanism. Since the covariates are not observed directly, we also need to put a model on it, for example $X_{i}\overset{iid}{\sim}f$.
Privatized Bayesian linear regression refers to inferring the posterior distribution of parameters $\theta=(\beta_{0},\beta,\sigma^{2})$ given privatized observations $s_{\mathrm{dp}}$. Data augmentation is a class of Gibbs samplers to make inference about model parameters in the presence of latent data. It entails an inference sample to update $\theta\mid X,y,s_{\mathrm{dp}}$ and an imputation step that updates $X,y\mid\theta,s_{\mathrm{dp}}$. The conditional distribution in the imputation step is permutation invariant and we can use SOMA in this step. When we use conjugate priors on $\theta$, the inference step is tractable. Convergence of the marginal $\theta$ chain is mainly determined by the convergence rate of the imputation step.
Method Coupling Time Convergence Rate Acceptance Rate (%)
$n=10,\ \epsilon=3$
Ran-IMwG $156.04\ (46,216)$ $0.9937\ (0.9933,0.9941)$ $94.80\ (94.52,95.11)$
Sys-IMwG $\textbf{99.60}\ (\textbf{28},\textbf{138})$ $\textbf{0.9904}\ (\textbf{0.9895},\textbf{0.9911})$ $94.81\ (94.55,95.09)$
SOMA $152.98\ (42,211)$ $0.9937\ (0.9932,0.9942)$ $\textbf{99.49}\ (\textbf{99.44},\textbf{99.55})$
$n=10,\ \epsilon=30$
Ran-IMwG $900.79\ (261,1262)$ $0.9989\ (0.9988,0.9990)$ $50.95\ (50.64,51.25)$
Sys-IMwG $549.07\ (170,764)$ $0.9982\ (0.9980,0.9983)$ $50.94\ (50.66,51.23)$
We first compare the coupling time of the joint posterior $(X,Y,\theta)\mid s_{\mathrm{dp}}$ for data augmentation MCMC using SOMA, Ran-IMwG, and Sys-IMwG for $n=10$. In Table 1, we also include the empirical acceptance rate of the data imputation steps. For a small privacy loss budget $\epsilon=3$, all three methods have empirical acceptance probability higher than 94% and they all converge fast. At $\epsilon=30$, the advantage of SOMA becomes evident: SOMA has an acceptance rate higher than $90\%$ while the other two samplers accept only $50\%$ of the proposed points. This translates into more rapid coupling and slightly faster convergence of the joint $(X,Y,\theta)$ chain. This joint distribution has dimension $n\cdot(p+2)+(p+1)=43.$
Next we compare the convergence of the marginal $\theta$ chain at $n=100.$ We no longer use coupling-based diagnosis as the dimension of the state space is $403$ at this point. We give a ridge plot of the intercept term $\beta_{0}$ in Figure 8 using $\epsilon=30$, showing how the posterior of $\beta$ given $s_{\mathrm{dp}}$ changes over MCMC iterations. Based on visual inspections, it takes a DAMCMC algorithm using SOMA about 150 iterations to get convergence in $\beta_{0}$, while Ran-IMwG and Sys-IMwG require 300 and 200 iterations, respectively.
A
B
@@ 5.3 Bayesian linear regression given privatized data
Figure 9: Convergence diagnosis using independent chains. A. R-hat diagnostic across iterations for three methods. The dashed line is 1.05. B. Effective sample size (ESS) across iterations for three methods
We use traditional convergence diagnosis such as the $\hat{R}$ (R-hat or the Gelman-Rubin diagnosis) and effective sample size (Aki et al.,, 2020). A large $\hat{R}$ indicates poorer mixing, while $\hat{R}\approx 1$ is generally considered indicative of good mixing. As shown in Figure 9-A, the chains utilizing SOMA require approximately 800 iterations to achieve $\hat{R}<1.05$, while both Sys-IMwG and Ran-IMwG require about 1,500 iterations. As shown in Figure 9-B, using SOMA within DAMCMC results in higher ESS than the other two methods.
Data availability statement
The code and data supporting this study are available from the author upon reasonable request.
Appendix A SOMA as a multi-proposal MCMC sampler

### v2

@@ 5 NUMERICAL EXPERIMENTS
In this section, we empirically compare SOMA with Sys-IMwG and Ran-IMwG samplers, in terms of convergence rates and acceptance probabilities, which have been studied in Section 4. We study a differentially private data release scenario in Section 5.1, and in Section 5.2, we embed SOMA into a data augmentation MCMC framework for private Bayesian linear regression, highlighting SOMA’s versatility. Moreover, Appendix B.1 provides an example of data imputation based on perturbed histograms.
5.1 Imputation Given a Privatized Mean
We first consider a canonical privacy problem: releasing a privatized mean of confidential data drawn from a bounded domain. Assuming $X_{i}\in[0,1]$ for each $i$, with model them by $X_{i}\sim\text{Beta}(a_{0},b_{0})$ independently with $n=2$. Let $Y\sim\textrm{Laplace}(({X_{1}+X_{2}})/{2},{1}/{\epsilon})$, which satisfy $\epsilon$-DP, be the perturbed mean. Given $Y=y$, the posterior density of the confidential data is given by
$\pi(x\mid y)\propto\$$\exp\left(-\frac{\epsilon}{2}\left|y-\frac{x_{1}+x_{2}}{2}\right|\right)\cdot(x_{1}x_{2})^{a_{0}-{}1}(1-x_{1})^{b_{0}-1}(1-x_{2})^{b_{0}-1}.$
@@ 5.1 Imputation Given a Privatized Mean
Figure 1 gives the 2-dimensional trace plots of the three methods over 500 iterations, starting from the same initial state $(0.3,0.3)$ with $a_{0}=b_{0}=10$ and $\epsilon=20$. SOMA traverses the higher density regions of $\pi$ more rapidly and can explore the state space more efficiently. In contrast, Ran-IMwG gets stuck in some local regions, resulting in an under-exploration of the entire state space.
@@
Figure 1: Trace plots of different methods. The starting point is marked in deep purple, with subsequent samples in lighter colors. The orange density curve represents the true marginal posterior. SOMA achieves better exploration compared to others.
Next, we quantitatively compare the methods using their empirical acceptance probabilities in Figure 2. SOMA has a higher acceptance rate than Ran-IMwG and SYS-IMwG, agreeing with the results proved in Section 4.
Figure 2: Average acceptance rate comparison for varying values of $\epsilon$.
Finally, in challenging scenarios with a strong correlation between $X_{1}$ and $X_{2}$, SOMA also exhibits the fastest convergence rate; See Appendix B.2 for more results.
5.2 Bayesian Linear Regression Given Privatized Data
This section concerns the full data augmentation MCMC algorithm targeting the joint distribution $\pi(\theta,x\mid s_{\mathrm{dp}})$ of (2) for linear regression.
The mechanism for releasing perturbed summaries has been described in Example 2. We will use conjugate priors for $\theta$, which makes the inference step described in Section 2.2 tractable. In this case, the convergence of the marginal $\theta$ chain is dictated mainly by the convergence rate of the imputation step. Appendix B.3 describes the data generating parameters for simulating the dataset.
First, consider the small sample size ($n=10$) case. The dimension of the joint chain is $n\cdot(p+2)+(p+1)=43.$ For a small privacy loss budget like $\epsilon=3$, SOMA’s acceptance rate is 99.49%, making it nearly rejection-free, whereas IMwG methods accept around 94.80%. SOMA’s advantages become more evident under a larger privacy budget ($\epsilon=30$), where it achieves an acceptance rate about 91.91% (compared to 50% for IMwG baselines) and couples significantly faster. Full results are available in Appendix 3.
Next we compare the convergence of the marginal $\theta$ chain at $n=100.$ We no longer use coupling-based diagnosis as the dimension of the state space is $n\cdot(p+2)+(p+1)=403$ at this point. We use traditional convergence diagnosis such as the $\hat{R}$ (R-hat or the Gelman-Rubin diagnosis) and effective sample size (Aki et al.,, 2020, ESS). A large $\hat{R}$ indicates poorer mixing, while $\hat{R}\approx 1$ is generally considered indicative of good mixing. As shown in Figure 3-A, the chains utilizing SOMA require approximately 800 iterations to achieve $\hat{R}<1.05$, while both Sys-IMwG and Ran-IMwG require about 1,500 iterations. As shown in Figure 3-B, using SOMA within DAMCMC results in a higher ESS than the other two methods.
@@ 5.2 Bayesian Linear Regression Given Privatized Data
Figure 3: Convergence diagnosis using independent chains. A. $\hat{R}$ diagnostic across iterations for three methods. The dashed line is 1.05. B. ESS across iterations for three methods
6 CENSUS DATA EXPERIMENT
We apply SOMA to the privacy-preserving analysis of compositional data using the 2019 ATUS microdata file, which is collected by the U.S. Bureau of Labor Statistics. The data is available at https://www.bls.gov/tus/datafiles-2019.htm. The confidential data $\{x_{i}\}_{i=1}^{n}$ represent the fraction of time spent on personal care ($x_{i1}$), eating and drinking ($x_{i2}$), and all other activities ($x_{i3}$). The privacy mechanism has been reviewed in Example 3. We use a privacy budget $\epsilon=10$ and a clipping lower bound $a=0.0006$.
The data is split by female and male. After data pre-processing, the sample consists of 3,528 females and 3,128 males. We model the data with a Dirichlet distribution with parameter vector $\alpha=(\alpha_{1},\alpha_{2},\alpha_{3})$ and specify a $\mathrm{Gamma}(1.0,0.1)$ prior for each $\alpha_{j}$, $j=1,2,3$. The same model, budget, clipping lower bound, and priors have also been used in Guo et al., (2024); Awan et al., (2024),
Unlike in Section 5.2, in this problem, the inference step $p(\alpha\mid x_{1},\ldots,x_{n})$ cannot be sampled exactly, and we resort to using a slice sampler (Neal,, 2003) to sample from it approximately. We compare using SOMA or IMwG for the imputation step.
Table 1 presents the chain diagnostics, comparing using SOMA versus using IMwG for the imputation step. We report the $\hat{R}$ after $10^{5}$ iterations and the ESS after $10^{6}$ iterations based on 20 independent runs, which shows that the SOMA-based update achieves a lower $\hat{R}$ and higher ESS across all dimensions.
Table 1: Diagnostic Summary for SOMA and IMwG
Method Female Male
$\hat{R}$ ESS $\hat{R}$ ESS
$\alpha_{1}$ SOMA 1.0469 2754.5 1.0492 3406.8
IMwG 1.1041 2373.5 1.1994 2249.0
$\alpha_{2}$ SOMA 1.0333 3576.5 1.0344 4510.3
IMwG 1.0816 2965.0 1.1546 2777.4
$\alpha_{3}$ SOMA 1.0481 2713.6 1.0506 3342.5
IMwG 1.1068 2348.7 1.2043 2216.6
Table 2 presents the estimated mean time allocations, $E[X\mid s_{\mathrm{dp}}]$. We use these posterior estimates to test for significant gender-based differences, evaluating the null hypothesis that the absolute difference in expected time allocation for each activity $j$ is no more than one percentage point ($H_{0,j}:|E[X_{j}|\text{Female}]-E[X_{j}|\text{Male}]|\leq 0.01$). Our analysis leads to the rejection of the null hypotheses for personal care ($H_{0,1}$) and other activities ($H_{0,3}$), while the null hypothesis for eating and drinking ($H_{0,2}$) is not rejected. Both the point estimates and hypothesis testing conclusions are consistent with the prior findings in Guo et al., (2024).
Table 2: Estimated Mean Time Allocation $E[X\mid s_{\mathrm{dp}}]$
Activity Female Male
SOMA IMwG SOMA IMwG
Personal Care 0.4112 0.4112 0.3919 0.3919
Eating & drinking 0.0507 0.0506 0.0508 0.0508
Other activities 0.5381 0.5382 0.5573 0.5574
Appendix A PROOFS AND TECHNICAL RESULTS
A.1 Proofs of Section 3
Proof of Theorem 5.
To verify the detailed balance condition, we should derive the ‘backward proposal kernel’ $K(x\mid x^{\prime})$. In this proposal kernel, the current state is $x^{\prime}$, the proposed point is $x_{1}$, the index $i$ is selected according to weights $w_{i}(x_{1},x^{\prime})$ defined in (7) & (9). Notice that from $x^{\prime}$, we can only move to $x^{\prime}_{-1},x_{1}]=x$ or $x^{\prime}_{-i},x_{1}]$ which is equivalent to $x_{-i},y]$ for $i\geq 2$ up to shuffling indices, in terms of having equal density under $\pi$. Using our simplified weight definitions, we have
$w_{1}(x_{1},x^{\prime})$ $=\eta(s_{\mathrm{dp}}\mid s([x^{\prime}_{-1},x_{1}]))=\eta(s_{\mathrm{dp}}\mid s(x))=w_{0}(y,x),$
$w_{k}(x_{1},x^{\prime})$ $=\eta(s_{\mathrm{dp}}\mid s([x^{\prime}_{-k},x_{1}]))=\eta(s_{\mathrm{dp}}\mid s([x_{-k},y]))=w_{k}(y,x),\quad\text{for }k\geq 2.$
Using these weights, the backward proposal kernel density is
$K(x\mid x^{\prime})=f(x_{1}\mid\theta)\frac{w_{1}(x_{1},x^{\prime})}{\sum_{k=1}^{n}w_{k}(x_{1},x^{\prime})}=f(x_{1}\mid\theta)\frac{w_{0}(y,x)}{w_{0}(y,x)+\sum_{k=2}^{n}w_{k}(y,x)}.$
We also have
$\frac{\pi(x^{\prime})K(x\mid x^{\prime})}{\pi(x)K(x^{\prime}\mid x)}=\frac{\pi(x^{\prime})f(x_{1}\mid\theta)\frac{w_{0}(y,x)}{w_{0}+\sum_{j=2}^{n}w_{j}(y,x)}}{\pi(x)f(y\mid\theta)\frac{w_{1}}{\sum_{j=1}^{n}w_{j}(y,x)}}=\frac{\sum_{j=1}^{n}w_{j}(y,x)}{w_{0}+\sum_{j=2}^{n}w_{j}(y,x)},$ (20)
now the following identity holds:
$\pi(x^{\prime})f(x_{1}\mid\theta)w_{0}(y,x)$ $=\left(\eta(s_{\mathrm{dp}}\mid s(x^{\prime}))\prod_{j\neq 1}f(x_{j}\mid\theta)f(y\mid\theta)\right)f(x_{1}\mid\theta)\eta(s_{\mathrm{dp}}\mid s(x))$
$=\left(\eta(s_{\mathrm{dp}}\mid s(x))\prod_{j\neq 1}f(x_{j}\mid\theta)f(x_{1}\mid\theta)\right)f(y\mid\theta)\eta(s_{\mathrm{dp}}\mid s(x^{\prime}))$
$=\pi(x)f(y\mid\theta)w_{1}(y,x).$
Proof of Corollary 6.
For the general cases, assume $x$ and $x^{\prime}$ differ only on component 1, i.e., $x^{\prime}=[x_{-1},y]$. The weights are now given by
$w_{1}(x_{1},x^{\prime})$ $=c\cdot\frac{\pi([x^{\prime}_{-1},x_{1}])}{q(x_{1})\prod_{j=2}^{n}q(x_{j}^{\prime})}=c\cdot\frac{\pi(x)}{\prod_{j=1}^{n}q(x_{j})}=w_{0}(y,x),$
$w_{k}(x_{1},x^{\prime})$ $=c\cdot\frac{\pi([x^{\prime}_{-k},x_{1}])}{q(x_{1})q(y)\prod_{2\leq j\leq n,j\not=k}q(x^{\prime}_{j})}=c\cdot\frac{\pi([x_{-k},y])}{q(y)\prod_{1\leq j\leq n,j\not=k}q(x_{j})}=w_{k}(y,x),\quad\text{for }k\geq 2,$
and we have the corresponding backward proposal kernel as
$K(x\mid x^{\prime})=q(x_{1})\frac{w_{1}(x_{1},x^{\prime})}{\sum_{k=1}^{n}w_{k}(x_{1},x^{\prime})}=q(x_{1})\frac{w_{0}(y,x)}{w_{0}(y,x)+\sum_{k=2}^{n}w_{k}(y,x)},$
which yields
$\frac{\pi(x^{\prime})K(x\mid x^{\prime})}{\pi(x)K(x^{\prime}\mid x)}=\frac{\pi(x^{\prime})q(x_{1})\frac{w_{0}(y,x)}{w_{0}+\sum_{j=2}^{n}w_{j}(y,x)}}{\pi(x)q(y)\frac{w_{1}}{\sum_{j=1}^{n}w_{j}(y,x)}}=\frac{\sum_{j=1}^{n}w_{j}(y,x)}{w_{0}+\sum_{j=2}^{n}w_{j}(y,x)}.$
Now since
$\pi(x^{\prime})q(x_{1})w_{0}(y,x)=\pi(x^{\prime})q(x_{1})\frac{c\pi(x)}{\prod_{j=1}^{n}q(x_{j})}=c\pi(x)\frac{q(y)}{q(y)}\frac{\pi([x_{-1},y])}{\prod_{j=2}^{n}q(x_{j})}=\pi(x)q(y)w_{1}(y,x),$
thus, the Metropolis-Hastings acceptance ratio $\alpha(x^{\prime}\mid x)=\min\left(1,\frac{\pi(x^{\prime})K(x\mid x^{\prime})}{\pi(x)K(x^{\prime}\mid x)}\right)$ still can yields (8), which makes the Markov transition kernel reversible with respect to $\pi$. ∎
A.2 Generalization to SOMA with $m<n$ components
We now provide a detailed discussion on the properties of the SOMA sampler when adapted to operate on a subset $M\subseteq\{1,\dots,n\}$ of size $m$.
First, when $m=1$, SOMA becomes deterministic or random scan, depending on how the index set is chosen.
Second, each SOMA kernel using $m$ components still satisfies detailed balance.
Lemma 14.
Use same conditions as Corollary 6. Let $P(x^{\prime}\mid x,M)$ be the transition probability for a given index set $M$. The SOMA sampler operating on $M$ satisfies the detailed balance condition with respect to $\pi(x)$:
$\pi(x)P(x^{\prime}\mid x,M)=\pi(x^{\prime})P(x\mid x^{\prime},M).$ (21)
Proof.
We verify detailed balance for $x\not=x^{\prime}$ and this is to check $\pi(x)K(x^{\prime}\mid x,M)\alpha(x^{\prime}\mid x,M)=\pi(x^{\prime})K(x\mid x^{\prime},M)\alpha(x\mid x^{\prime}.M)$ Let $i$ be the index where $x_{i}\not=x^{\prime}_{i}$. If $i\not\in M$, then both sides of (21) are zero. The proposal density for moving from $x$ to $x^{\prime}$ (by proposing $y=x^{\prime}_{i}$ and selecting index $i\in M$) has the form
[…]
and the backward move from $x^{\prime}$ to $x$ involves proposing the $x_{i}$ to replace $x_{i}^{\prime}$, which yields
[…]
Recall also that
$\alpha(x^{\prime}\mid x,M)$ $=\min\left(1,\frac{W_{M}(x^{\prime}_{i},x)}{w_{0}(x^{\prime}_{i},x)+\sum_{j\not=i,j\in M}w_{j}(x^{\prime}_{i},x)}\right)=\min\left(1,\frac{W_{M}(x^{\prime}_{i},x)}{w_{0}(x^{\prime}_{i},x)+W_{M}(x^{\prime}_{i},x)-w_{i}(x^{\prime}_{i},x)}\right),$
By construction of the weights (7) and (9), we have $w_{0}(x^{\prime}_{i},x)=w_{i}(x_{i},x^{\prime})$ and $w_{0}(x_{i},x^{\prime})=w_{i}(x^{\prime}_{i},x).$ More importantly $w_{j}(x_{i}^{\prime},x)=w_{j}(x_{i},x^{\prime})$ for each $j\in M$ due to permutation invariance. Now we have
$W^{\prime}_{M}(x_{i},x^{\prime})$ $=w_{i}(x_{i},x^{\prime})+\sum_{j\not=i,j\in M}w_{j}(x_{i},x^{\prime})$
$=w_{0}(x^{\prime}_{i},x)+\sum_{j\not=i,j\in M}w_{j}(x^{\prime}_{i},x)$
$=w_{0}(x^{\prime}_{i},x)-w_{i}(x^{\prime}_{i},x)+W_{M}(x^{\prime}_{i},x).$
We can check that
$\frac{\pi(x^{\prime})K(x\mid x^{\prime},M)}{\pi(x)K(x^{\prime}\mid x,M)}$ $=\frac{\pi(x^{\prime})\cdot q(x_{i})\frac{w_{i}(x_{i},x^{\prime})}{W^{\prime}_{M}(x_{i},x^{\prime})}}{\pi(x)\cdot q(x^{\prime}_{i})\frac{w_{i}(x^{\prime}_{i},x)}{W_{M}(x^{\prime}_{i},x)}}=\frac{\pi(x^{\prime})\cdot q(x_{i})w_{i}(x_{i},x^{\prime})W_{M}(x^{\prime}_{i},x)}{\pi(x)\cdot q(x^{\prime}_{i})w_{i}(x^{\prime}_{i},x)W^{\prime}_{M}(x_{i},x^{\prime})}$
$=\frac{\pi(x^{\prime})q(x_{i})\cdot w_{0}(x^{\prime}_{i},x)}{\pi(x)q(x^{\prime}_{i})\cdot w_{0}(x_{i},x^{\prime})}\cdot\frac{W_{M}(x^{\prime}_{i},x)}{w_{0}(x^{\prime}_{i})-w_{i}(x^{\prime}_{i},x)+W_{M}(x^{\prime}_{i},x)}$
$=\frac{W_{M}(x^{\prime}_{i},x)}{W_{M}(x^{\prime}_{i},x)+w_{0}(x^{\prime}_{i},x)-w_{i}(x^{\prime}_{i},x)}.$
Setting $\alpha(x^{\prime}\mid x,M)=\min\left(1,\frac{\pi(x^{\prime})K(x\mid x^{\prime},M)}{\pi(x)K(x^{\prime}\mid x,M)}\right)$ satisfies detailed balance and hence makes the Markov transition kernel reversible with respect to $\pi.$
Finally, SOMA remains reversible with respect to the target distribution $\pi(x)$ when using randomly selected subsets. Assume that at each iteration, a subset $M$ of size $m$ is drawn uniformly at random from all possible subsets of $\{1,\dots,n\}$ of that size. The overall transition kernel $P_{m}(x^{\prime}\mid x)$ becomes
$P_{m}(x^{\prime}\mid x)=\frac{1}{\binom{n}{m}}\sum_{M}P(x^{\prime}\mid x,M).$
With Lemma 14, the reversibility for the random subset case is straightforward by verifying the detailed balance equation as
$\pi(x)P_{m}(x^{\prime}\mid x)=\frac{1}{\binom{n}{m}}\pi(x)\sum_{M}P(x^{\prime}\mid x,M)=\frac{1}{\binom{n}{m}}\pi(x^{\prime})\sum_{M}P(x\mid x^{\prime},M)=\pi(x^{\prime})P_{m}(x\mid x^{\prime}).$
Thus, the SOMA sampler with randomly chosen subsets is still reversible with respect to the target distribution $\pi(x)$.
A.3 SOMA as a multi-proposal sampler